OpenStudy (anonymous):

temperature in C on Feb 5 and 6 can be approximated using T(t)=20=5cos(pi/12(t-3)) where t is time in hours after noon on Feb 5 What were the min and max temps and what time did they occur. Could I get some help please

OpenStudy (perl):

the min of a cosine function occurs when the argument is pi

OpenStudy (perl):

The minimum occurs at T = 20 + 5cos( Pi ) when pi/12 (t-3) = Pi

OpenStudy (anonymous):

so 1500 hours?

OpenStudy (anonymous):

help please :)

OpenStudy (perl):

Did you solve pi/12 (t - 3) = pi (t-3) = 12 t = 15